Elise and Joe purchased a new combine for $56,000. Assume a 7 year useful life and no salvage value. Calculate the depreciation for each year of the useful life using: A. straight-line methoo B. double declining-balance method C. sum-of-the-years-digits method D. MACRS (assume 7 year property and midyear convention Double Declining- Sum-of-the- Years-Digits Stright-Line Balance MACRS Year 1 Year 2 Year 3 Year 4 Year 5 Year 6 Year 7 Year 8
Answer #1

A. Straight Line Method

Depreciation charges = $ 56,000/7 = $ 8000 every year

B. Double Declining Balance

Depreciation charges = $56000*2/7 = $ 16,000

Second year = $(56000-16000)*2/7 = 11428.57

Like wise we can calculate.

In case of sum of years digit add the digit of the years

1+2+3+4+5+6+7 = 28

Depreciation charges of 1st year = $ 56000*7/28 = $ 14,000

Second year = $ 56,000*6/28 = $ 12,000

3rd year = $56000*5/28 = $ 10000

Likewise we can calculate for other years too.

MACRS rate for 7 year property is

1st year 14.29%

2nd year 17.49%

3rd year 24.49%

4th year 12.49%

5th year 8.93%

6th year 8.92%

7th year 8.93%

8th year 4.46%

Multiply these rates with $ 56000 you will get depreciation charges.

Please refer the attached picture

Straight line| Double Declining |Sum of years digitMacrs 8002.4 13714.4 9794.4 6994.4 5000.8 4995.2 5000.8 2497.6 8000 8000 8000 8000 8000 8000 8000 16000 11428.57 8163.27 5830.90 4164.93 2974.95 2124.96 14000 12000 10000 8000 6000 4000 2000 1 6

Please contact if having any query thank you.